# Shezheng Song*, Chengxiang He*, Shasha Li, Shan Zhao, Chengyu Wang Tianwei Yan, Xiaopeng Li, Qian Wan, Jun Ma, Jie Yu, Xiaoguang Mao

### v1

@@ Abstract
Index Terms:
Multimodal Large Language Model, Multimodal Sentiment Analysis, Benchmark.
publicationid: pubid: $0$–$00$ 2021 IEEE
@@ I Introduction
In recent years, multimodal large language models [1] (MLLMs) have made significant progress in image understanding tasks, demonstrating immense potential, particularly in high-level semantic tasks such as visual question answering [2, 3, 4], image captioning [5, 6], and emotion recognition [7, 8, 9]. Commercial platforms like GPT4o [10] and Gemini [11], as well as open-source models such as Flamingo [12] and LLaVA [13], have achieved outstanding performance on a wide range of traditional tasks, even surpassing human-level performance in tasks like multimodal named entity recognition [14, 15, 16]. This impressive potential has driven the development of multimodal evaluation tasks. However, sentiment analysis [17, 18], as one of the key tasks in semantic understanding, still lacks a standardized benchmark specifically designed for evaluating MLLMs. Figure 1 shows that the performance of MLLMs on our MOSABench is unsatisfactory.
As shown in Figure 2 and 3, current sentiment analysis datasets have limitations in accurately evaluating the capabilities of MLLMs. Lack of focus on multi-object understanding: In real-world social media, images typically contain multiple objects (such as people), each of which may convey different emotional information. Therefore, MLLMs are facing the challenge of multiple object sentiment analysis. However, most existing sentiment analysis benchmarks, such as Twitter15, Twitter17, and MSED [19], are dominated by single-object samples, which leads to models focusing primarily on single-object sentiment classification ability [20, 21]. When single-object data dominate the dataset, the model’s good performance likely reflects its ability to classify emotions for individual objects, rather than its capacity to analyze emotions across multiple objects in an image. As a result, evaluating models on such imbalanced datasets fails to comprehensively assess their ability to understand multiple-object emotions, which limits the potential of multimodal sentiment analysis models. Lack of adaptability to MLLMs: Moreover, existing datasets [22, 19] have substantial limitations in evaluating the capabilities of MLLMs. Most of these datasets are designed for smaller models, lack instruction, and require extensive adaptation, making it challenging to comprehensively assess the performance of MLLMs. Additionally, since MLLMs generate outputs in varying formats rather than strictly following predefined formats, accurately evaluating their performance is more difficult. Therefore, developing a standardized, multidimensional benchmark specifically designed for evaluating multi-object sentiment analysis tasks has become an urgent challenge.
@@
We conduct a comprehensive evaluation of mainstream MLLMs on our MOSABench. Our analysis leads to the following key observations: (1) Challenges in MLLMs for multiple objects analysis: Most current models perform suboptimally on the MOSABench multi-object sentiment analysis task, revealing significant challenges in handling multiple target emotions simultaneously. Although these models perform well on other tasks, such as VQA [2, 23] and MNER [14, 15], they have limitations when it comes to multi-object sentiment analysis, which demands higher comprehension and reasoning abilities for multiple objects. (2) Impact of object distance on accuracy: Our further analysis reveals a significant correlation between task accuracy and the spatial relationship between targets in the image. We categorize the distance between objects in the data into three types: Interlap, Close, and Far. The experimental results show that as the distance between targets increases, the performance of most MLLMs decreases. This finding provides a new perspective for improving model performance in multi-object sentiment analysis and emphasizes the importance of spatial relationships in multi-object contexts. (3) Performance variations across models: We also observe significant performance differences across models. For example, the mPLUG-owl [24] and Qwen-VL2 [25] models demonstrate relatively high stability and accuracy across various evaluation metrics. This may be due to their optimized architectures and larger parameter scales, which enable them to better adapt to complex sentiment analysis tasks. In contrast, models such as Qwen-VL and BLIVA-Flant5 show poorer performance, suggesting a lack of sufficient generalization capability in multi-object sentiment analysis, making it difficult for these models to accurately capture emotional features of multiple objects.
@@ II-A Advances in MLLMs for Sentiment Analysis and the Demand for Scientific Benchmarking
The recent advancements in large language models (LLMs) [26, 27, 28, 29] have significantly improved multimodal sentiment analysis, effectively handling complex sentiment tasks across various modalities [30, 31]. Studies have demonstrated that LLMs achieve high accuracy on standard datasets, such as Twitter15 and Twitter17 [22], which are widely used to assess sentiment analysis capabilities by integrating text and image data to analyze social media posts. For example, the WisdoM[32] framework leverages large vision-language models (LVLMs) to enhance sentiment analysis by incorporating contextual world knowledge from images and text, thus improving LLMs interpretability and performance [33]. Similarly, the PSL [34] framework is a pipeline-based approach for aspect-based sentiment analysis, using small language models (SLMs) [35] for aspect extraction and MLLMs for sentiment analysis. This structured guidance enables MLLMs to focus on relevant image regions, effectively addressing the complexities of multimodal sentiment tasks [34]. These frameworks highlight the progress LLMs have made in multimodal alignment and sentiment understanding [20, 33]. However, current benchmarks, like Twitter15 and Twitter17 [22], reveal limitations in assessing MLLMs true multimodal comprehension capabilities. Primarily, these datasets often lack image-text consistency, where the targets mentioned in the text may not appear in the associated image, hindering accurate evaluations of MLLMs capacity to integrate visual context [33]. Additionally, these benchmarks are not equipped with LLM-specific instructions [36], making it challenging to assess the impact of different prompting methods on sentiment prediction [34]. Lastly, they lack multi-object sentiment assessment, which is crucial for evaluating the ability of MLLMs to independently analyze sentiments towards multiple entities within a single post [34, 32]. These gaps underscore the need for a scientifically designed benchmark that captures multimodal nuances, includes structured prompts tailored for LLM tasks, and offers multi-object sentiment assessment to comprehensively evaluate LLM performance in real-world multimodal sentiment applications.
@@ II-B Limitations of Current Benchmarks in Addressing Sentiment Analysis Needs
In the evaluation of LLMs, tasks such as Named Entity Recognition (MNER) [15, 14] have developed dedicated benchmarks to better measure model performance. However, sentiment analysis lacks a specialized benchmark tailored specifically for LLMs. While some comprehensive benchmarks, such as MM-SOC [37] and MM-BigBench [38], include multimodal tasks related to sentiment analysis, they still exhibit notable limitations in design and evaluation [39, 40], particularly in areas of image-text consistency, instruction design, and multi-object sentiment judgment [32, 34]. For instance, MM-SOC primarily targets multimodal tasks within social media environments, covering sentiment analysis, hate speech detection [41], and more, using the Memotion dataset [42] for joint image-text emotion detection. However, MM-SOC does not specifically address image-text consistency, which limits the evaluation of whether all entities mentioned in the text are also represented in the image. Additionally, MM-SOC lacks LLM-specific instructions, restricting its utility in large-scale multimodal tasks that require instruction-following capabilities [34]. MM-BigBench, on the other hand, covers a broad range of multimodal comprehension tasks, including visual question answering and multimodal sentiment analysis (MSA), focusing on multimodal information fusion and deep understanding. However, MM-BigBench does not provide detailed evaluations for multi-object sentiment analysis and does not adequately emphasize image-text consistency, which is essential for accurately identifying the sentiments toward multiple entities mentioned in the text. Furthermore, this benchmark lacks instructions specifically designed for LLMs, making it difficult to assess how different prompt structures might affect performance.
@@ III-A Dataset Construction
4:    $C_{1}=\left(\frac{x_{1}+x_{1}^{\prime}}{2},\frac{y_{1}+y_{1}^{\prime}}{2}\right)$
5:    $C_{2}=\left(\frac{x_{2}+x_{2}^{\prime}}{2},\frac{y_{2}+y_{2}^{\prime}}{2}\right)$triangleright Calculate the midpoint of each bounding box
@@ III-C Datasets Statistics
The sentiment distribution in MOSABench reflects a scientifically consistent approach, aligning with the proportions observed in prior datasets. In MOSABench, Negative, Neutral, and Positive sentiments are represented by 15.44%, 49.43%, and 35.13%, respectively. This design aligns closely with the sentiment distributions in the Twitter15 and Twitter17 datasets [22], which feature similar proportions: Twitter15 includes 12.06% Negative, 59.29% Neutral, and 28.65% Positive samples, while Twitter17 comprises 12.19% Negative, 45.68% Neutral, and 42.13% Positive samples. By mirroring these established distributions, MOSABench provides a scientifically robust benchmark for evaluating MLLMs sentiment analysis capabilities across multiple targets.
@@ III-D MLLM Baselines
Based on the MOSABench dataset we constructed, we conduct a comprehensive evaluation of existing MLLMs to assess their performance in multi-object sentiment analysis tasks. To this end, we propose two baseline methods and systematically test various mainstream model architectures, covering three distinct types. In terms of model selection, we choose representative MLLMs, categorized into three groups according to their architectures:
  • Open-sourced models: LLaVA1.6 [13], mPLUG-Owl [24], Qwen-VL [43], Qwen-VL2 [25], VisualGLM [44], BLIVA-FlanT5 [45], Monkey [46], GLM4V [47], InternLM2.5 [48].
  • Close-sourced models: ERNIE-Bot [49], GPT-4o [10] and Gemini [11].
@@ III-F Metrics
LLaVA1.6-7B[13] 51.31 51.31 46.87 53.92 59.08 49.58 51.30 57.30 46.43 44.26 47.37 41.54
mPLUG-owl-7B[24] 73.16 69.86 76.78 72.80 70.18 75.62 74.08 70.66 77.85 69.53 65.10 74.62
Qwen-VL-7B[43] 33.81 29.14 40.26 36.98 31.86 44.04 33.55 28.93 39.91 26.37 22.65 31.54
Qwen-VL2-7B[25] 58.39 61.72 55.39 60.53 64.09 57.34 58.17 61.63 55.08 53.60 55.83 51.54
VisualGLM-6B[44] 16.71 14.22 20.26 15.97 13.72 19.11 17.59 14.94 21.40 14.33 12.04 17.69
BLIVA-Flant5[50] 16.59 14.54 19.30 15.82 20.78 17.96 15.90 13.93 18.51 16.29 14.12 19.23
Monkey[46] 39.87 33.63 48.96 40.77 34.48 49.86 40.64 34.35 49.77 33.64 27.92 42.31
GLM4V-9B[51] 54.48 57.10 52.09 53.94 56.92 51.25 55.62 58.56 52.96 50.39 50.78 50.00
InternLM2.5-7B[48] 50.73 53.36 48.35 51.63 52.91 50.42 52.05 55.29 49.17 41.32 44.64 38.46
@@
ERNIE Bot[49] 68.62 70.18 67.13 69.79 71.51 68.14 69.05 70.81 67.37 63.32 63.57 63.08
GPT4o[10] 48.10 48.92 47.30 50.71 51.88 49.58 47.84 48.67 47.04 42.31 42.31 42.31
Gemini[11] 57.26 60.29 54.52 57.97 60.79 55.40 57.92 61.25 54.93 52.00 54.17 50.00
@@
LLaVA1.6-7B[13] 29.39 32.86 37.75 1.21 31.80 33.33 34.87 1.29 29.29 33.47 37.24 1.21 23.47 28.57 47.96 0.99
mPLUG-owl-7B[24] 68.70 20.07 11.23 2.26 67.05 19.54 13.41 2.21 69.67 20.50 9.83 2.29 68.37 19.39 12.24 2.24
Qwen-VL-7B[43] 26.16 27.12 46.71 1.06 28.74 30.27 41.00 1.16 26.36 26.78 46.86 1.06 18.37 20.41 61.22 0.76
Qwen-VL2-7B[25] 40.86 31.18 27.96 1.54 42.15 33.33 24.52 1.60 40.79 31.59 27.62 1.54 37.76 23.47 38.78 1.37
VisualGLM-6B[44] 15.17 8.60 76.22 0.54 15.71 7.66 76.63 0.55 15.27 9.41 75.31 0.55 13.27 7.14 79.59 0.47
BLIVA-Flant5[50] 14.10 12.43 73.48 0.55 14.56 14.18 71.26 0.58 14.85 10.67 74.48 0.55 9.18 16.33 74.49 0.44
Monkey[46] 33.57 28.55 37.87 1.29 35.25 29.50 35.25 1.35 34.10 28.87 37.03 1.31 26.53 24.49 48.98 1.04
GLM4V-9B[51] 36.44 31.30 32.26 1.41 34.48 33.72 31.80 1.37 37.66 31.59 30.75 1.45 35.71 23.47 40.82 1.31
InternLM2.5-7B[48] 32.97 30.47 36.56 1.29 34.10 31.42 34.48 1.34 34.52 30.54 34.94 1.33 22.45 27.55 50.00 0.95
@@
ERNIE Bot[49] 55.91 24.85 19.24 1.93 55.94 27.20 16.86 1.95 56.49 25.73 17.78 1.95 53.06 14.29 32.65 1.73
GPT4o[10] 34.05 26.05 39.90 1.28 35.63 27.97 36.40 1.35 34.10 26.15 39.75 1.28 29.59 20.41 50.00 1.09
Gemini[11] 39.19 30.59 30.23 1.48 39.46 32.57 27.97 1.51 39.96 31.17 28.87 1.51 32.69 22.45 42.86 1.27

### v2

# Shezheng Song, Chengxiang He, Shan Zhao, Chengyu Wang Qian Wan, Tianwei Yan, Meng Wang, IEEE Fellow

@@ Abstract
publicationid: pubid: 0000–0000/00\$00.00 © 2021 IEEE
@@ I Introduction
In recent years, multimodal large language models [yin2024survey] (MLLMs) have made significant progress in image understanding tasks, demonstrating immense potential, particularly in high-level semantic tasks such as visual question answering [antol2015vqa, wu2017visual], image captioning [hossain2019comprehensive], and emotion recognition [dzedzickis2020human, TMMmethod1, TMMmethod2, TMMmethod3]. Commercial platforms like GPT4o [GPT4] and Gemini [geminiteam2024geminifamilyhighlycapable], as well as open-source models such as Flamingo [Flamingo] and LLaVA [llava], have achieved outstanding performance on a wide range of traditional tasks, even surpassing human-level performance in tasks like multimodal named entity recognition [jia2023mner, ji2024cmner, TCSVT_ERC]. This impressive potential has driven the development of multimodal evaluation tasks. However, sentiment analysis [das2023multimodal, kaur2022multimodal], as one of the key tasks in semantic understanding, still lacks a standardized benchmark specifically designed for evaluating MLLMs. Figure 1 shows that the performance of MLLMs on our MOSABench is unsatisfactory.
As shown in Figure 2 and 3, current sentiment analysis datasets have limitations in accurately evaluating the capabilities of MLLMs. Lack of focus on multi-object understanding: In real-world social media, images typically contain multiple objects (such as people), each of which may convey different emotional information. Therefore, MLLMs are facing the challenge of multiple object sentiment analysis. However, most existing sentiment analysis benchmarks, such as Twitter15, Twitter17, and MSED [jia2022beyond], are dominated by single-object samples, which leads to models focusing primarily on single-object sentiment classification ability [phan2019sentiment, TCSVT2023mcl]. When single-object data dominate the dataset, the model’s good performance likely reflects its ability to classify emotions for individual objects, rather than its capacity to analyze emotions across multiple objects in an image. As a result, evaluating models on such imbalanced datasets fails to comprehensively assess their ability to understand multiple-object emotions, which limits the potential of multimodal sentiment analysis models. Lack of adaptability to MLLMs: Moreover, existing datasets [twitterdataset, jia2022beyond] have substantial limitations in evaluating the capabilities of MLLMs. Most of these datasets are designed for smaller models, lack instruction, and require extensive adaptation, making it challenging to comprehensively assess the performance of MLLMs. Additionally, since MLLMs generate outputs in varying formats rather than strictly following predefined formats, accurately evaluating their performance is more difficult. Therefore, developing a standardized, multidimensional benchmark specifically designed for evaluating multi-object sentiment analysis tasks has become an urgent challenge.
@@
We conduct a comprehensive evaluation of mainstream MLLMs on our MOSABench. Our analysis leads to the following key observations: (1) Challenges in MLLMs for multiple objects analysis: Most current models perform suboptimally on the MOSABench multi-object sentiment analysis task, revealing significant challenges in handling multiple target emotions simultaneously. Although these models perform well on other tasks, such as VQA [antol2015vqa, goyal2017VQA] and MNER [jia2023mner, ji2024cmner], they have limitations when it comes to multi-object sentiment analysis, which demands higher comprehension and reasoning abilities for multiple objects. (2) Impact of object distance on accuracy: Our further analysis reveals a significant correlation between task accuracy and the spatial relationship between targets in the image. We categorize the distance between objects in the data into three types: Interlap, Close, and Far. The experimental results show that as the distance between targets increases, the performance of most MLLMs decreases. This finding provides a new perspective for improving model performance in multi-object sentiment analysis and emphasizes the importance of spatial relationships in multi-object contexts. (3) Performance variations across models: We also observe significant performance differences across models. For example, the mPLUG-owl [mPLUG-Owl] and Qwen-VL2 [Qwen2VL] models demonstrate relatively high stability and accuracy across various evaluation metrics. This may be due to their optimized architectures and larger parameter scales, which enable them to better adapt to complex sentiment analysis tasks. In contrast, models such as Qwen-VL and BLIVA-Flant5 show poorer performance, suggesting a lack of sufficient generalization capability in multi-object sentiment analysis, making it difficult for these models to accurately capture emotional features of multiple objects.
@@ II-A Advances in MLLMs for Sentiment Analysis and the Demand for Scientific Benchmarking
The recent advancements in large language models (LLMs) [Minigpt-4, shikra, wei2023lenna, IdealGPT] have significantly improved multimodal sentiment analysis, effectively handling complex sentiment tasks across various modalities [li2023multimodal, MultiModal-GPT]. Studies have demonstrated that LLMs achieve high accuracy on standard datasets, such as Twitter15 and Twitter17 [twitterdataset], which are widely used to assess sentiment analysis capabilities by integrating text and image data to analyze social media posts. For example, the WisdoM[wang2024wisdom] framework leverages large vision-language models (LVLMs) to enhance sentiment analysis by incorporating contextual world knowledge from images and text, thus improving LLMs interpretability and performance [krugmann2024sentimentai]. Similarly, the PSL [psl] framework is a pipeline-based approach for aspect-based sentiment analysis, using small language models (SLMs) [DTCA] for aspect extraction and MLLMs for sentiment analysis. This structured guidance enables MLLMs to focus on relevant image regions, effectively addressing the complexities of multimodal sentiment tasks [psl]. These frameworks highlight the progress LLMs have made in multimodal alignment and sentiment understanding [phan2019sentiment, krugmann2024sentimentai]. However, current benchmarks, like Twitter15 and Twitter17 [twitterdataset], reveal limitations in assessing MLLMs true multimodal comprehension capabilities. Primarily, these datasets often lack image-text consistency, where the targets mentioned in the text may not appear in the associated image, hindering accurate evaluations of MLLMs capacity to integrate visual context [krugmann2024sentimentai]. Additionally, these benchmarks are not equipped with LLM-specific instructions [DreamLLM], making it challenging to assess the impact of different prompting methods on sentiment prediction [psl]. Lastly, they lack multi-object sentiment assessment, which is crucial for evaluating the ability of MLLMs to independently analyze sentiments towards multiple entities within a single post [psl, wang2024wisdom]. These gaps underscore the need for a scientifically designed benchmark that captures multimodal nuances, includes structured prompts tailored for LLM tasks, and offers multi-object sentiment assessment to comprehensively evaluate LLM performance in real-world multimodal sentiment applications.
@@ II-B Limitations of Current Benchmarks in Addressing Sentiment Analysis Needs
In the evaluation of LLMs, tasks such as Named Entity Recognition (MNER) [ji2024cmner, jia2023mner] have developed dedicated benchmarks to better measure model performance. However, sentiment analysis lacks a specialized benchmark tailored specifically for LLMs. While some comprehensive benchmarks, such as MM-SOC [jin2024mmsoc] and MM-BigBench [yang2023mmbench], include multimodal tasks related to sentiment analysis, they still exhibit notable limitations in design and evaluation [guo2023evaluating, wang2023evaluation], particularly in areas of image-text consistency, instruction design, and multi-object sentiment judgment [wang2024wisdom, psl]. For instance, MM-SOC primarily targets multimodal tasks within social media environments, covering sentiment analysis, hate speech detection [pi2023detgpt], and more, using the Memotion dataset [ramamoorthy2022memotion] for joint image-text emotion detection. However, MM-SOC does not specifically address image-text consistency, which limits the evaluation of whether all entities mentioned in the text are also represented in the image. Additionally, MM-SOC lacks LLM-specific instructions, restricting its utility in large-scale multimodal tasks that require instruction-following capabilities [psl]. MM-BigBench, on the other hand, covers a broad range of multimodal comprehension tasks, including visual question answering and multimodal sentiment analysis (MSA), focusing on multimodal information fusion and deep understanding. However, MM-BigBench does not provide detailed evaluations for multi-object sentiment analysis and does not adequately emphasize image-text consistency, which is essential for accurately identifying the sentiments toward multiple entities mentioned in the text. Furthermore, this benchmark lacks instructions specifically designed for LLMs, making it difficult to assess how different prompt structures might affect performance.
@@ III-A Dataset Construction
[…]
@@ III-C Datasets Statistics
The sentiment distribution in MOSABench reflects a scientifically consistent approach, aligning with the proportions observed in prior datasets. In MOSABench, Negative, Neutral, and Positive sentiments are represented by 15.44%, 49.43%, and 35.13%, respectively. This design aligns closely with the sentiment distributions in the Twitter15 and Twitter17 datasets [twitterdataset], which feature similar proportions: Twitter15 includes 12.06% Negative, 59.29% Neutral, and 28.65% Positive samples, while Twitter17 comprises 12.19% Negative, 45.68% Neutral, and 42.13% Positive samples. By mirroring these established distributions, MOSABench provides a scientifically robust benchmark for evaluating MLLMs sentiment analysis capabilities across multiple targets.
@@ III-D MLLM Baselines
Based on the MOSABench dataset we constructed, we conduct a comprehensive evaluation of existing MLLMs to assess their performance in multi-object sentiment analysis tasks. To this end, we propose two baseline methods and systematically test various mainstream model architectures, covering three distinct types. In terms of model selection, we choose representative MLLMs:
  • Open-sourced models: LLaVA1.6 [llava], mPLUG-Owl [mPLUG-Owl], Qwen-VL [qwen_vl], Qwen-VL2 [Qwen2VL], VisualGLM [du2022glm], BLIVA-FlanT5 [hu2023bliva], Monkey [li2024monkey], GLM4V [wang2023cogvlm], InternLM2.5 [InternLM2TechiReport].
  • Close-sourced models: ERNIE-Bot [sun2020ernie], GPT-4o [GPT4] and Gemini [geminiteam2024geminifamilyhighlycapable].
@@ III-F Metrics
LLaVA1.6-7B[llava] 51.31 51.31 46.87 53.92 59.08 49.58 51.30 57.30 46.43 44.26 47.37 41.54
mPLUG-owl-7B[mPLUG-Owl] 73.16 69.86 76.78 72.80 70.18 75.62 74.08 70.66 77.85 69.53 65.10 74.62
Qwen-VL-7B[qwen_vl] 33.81 29.14 40.26 36.98 31.86 44.04 33.55 28.93 39.91 26.37 22.65 31.54
Qwen-VL2-7B[Qwen2VL] 58.39 61.72 55.39 60.53 64.09 57.34 58.17 61.63 55.08 53.60 55.83 51.54
VisualGLM-6B[du2022glm] 16.71 14.22 20.26 15.97 13.72 19.11 17.59 14.94 21.40 14.33 12.04 17.69
BLIVA-Flant5[BLIVA] 16.59 14.54 19.30 15.82 20.78 17.96 15.90 13.93 18.51 16.29 14.12 19.23
Monkey[li2024monkey] 39.87 33.63 48.96 40.77 34.48 49.86 40.64 34.35 49.77 33.64 27.92 42.31
GLM4V-9B[CogVLM] 54.48 57.10 52.09 53.94 56.92 51.25 55.62 58.56 52.96 50.39 50.78 50.00
InternLM2.5-7B[InternLM2TechiReport] 50.73 53.36 48.35 51.63 52.91 50.42 52.05 55.29 49.17 41.32 44.64 38.46
@@
ERNIE Bot[sun2020ernie] 68.62 70.18 67.13 69.79 71.51 68.14 69.05 70.81 67.37 63.32 63.57 63.08
GPT4o[GPT4] 48.10 48.92 47.30 50.71 51.88 49.58 47.84 48.67 47.04 42.31 42.31 42.31
Gemini[geminiteam2024geminifamilyhighlycapable] 57.26 60.29 54.52 57.97 60.79 55.40 57.92 61.25 54.93 52.00 54.17 50.00
@@
LLaVA1.6-7B[llava] 29.39 32.86 37.75 1.21 31.80 33.33 34.87 1.29 29.29 33.47 37.24 1.21 23.47 28.57 47.96 0.99
mPLUG-owl-7B[mPLUG-Owl] 68.70 20.07 11.23 2.26 67.05 19.54 13.41 2.21 69.67 20.50 9.83 2.29 68.37 19.39 12.24 2.24
Qwen-VL-7B[qwen_vl] 26.16 27.12 46.71 1.06 28.74 30.27 41.00 1.16 26.36 26.78 46.86 1.06 18.37 20.41 61.22 0.76
Qwen-VL2-7B[Qwen2VL] 40.86 31.18 27.96 1.54 42.15 33.33 24.52 1.60 40.79 31.59 27.62 1.54 37.76 23.47 38.78 1.37
VisualGLM-6B[du2022glm] 15.17 8.60 76.22 0.54 15.71 7.66 76.63 0.55 15.27 9.41 75.31 0.55 13.27 7.14 79.59 0.47
BLIVA-Flant5[BLIVA] 14.10 12.43 73.48 0.55 14.56 14.18 71.26 0.58 14.85 10.67 74.48 0.55 9.18 16.33 74.49 0.44
Monkey[li2024monkey] 33.57 28.55 37.87 1.29 35.25 29.50 35.25 1.35 34.10 28.87 37.03 1.31 26.53 24.49 48.98 1.04
GLM4V-9B[CogVLM] 36.44 31.30 32.26 1.41 34.48 33.72 31.80 1.37 37.66 31.59 30.75 1.45 35.71 23.47 40.82 1.31
InternLM2.5-7B[InternLM2TechiReport] 32.97 30.47 36.56 1.29 34.10 31.42 34.48 1.34 34.52 30.54 34.94 1.33 22.45 27.55 50.00 0.95
@@
ERNIE Bot[sun2020ernie] 55.91 24.85 19.24 1.93 55.94 27.20 16.86 1.95 56.49 25.73 17.78 1.95 53.06 14.29 32.65 1.73
GPT4o[GPT4] 34.05 26.05 39.90 1.28 35.63 27.97 36.40 1.35 34.10 26.15 39.75 1.28 29.59 20.41 50.00 1.09
Gemini[geminiteam2024geminifamilyhighlycapable] 39.19 30.59 30.23 1.48 39.46 32.57 27.97 1.51 39.96 31.17 28.87 1.51 32.69 22.45 42.86 1.27